the outer faces of rectangular slab made up of equal thickness of iron and brass are maintained at 100 degree and 0 degree celcius given the thermal conductivities of iron and brass are 0.2 and 0.3 respectively. how to find the temperature at the interface?
 

Grade:11

1 Answers

GURIYA GUPTA
5 years ago
Rate of flow of heat energy through iron=rate of flow of heat energy through brass
KA(dT)/l= KA(dT)/l
As Area and length are same, therfore,
0.2(100-T)=0.3(T-0)
20-0.2T=0.3T
T=20/0.5
T=40°C